OpenStudy (anonymous):

find the derivative of... (x^2)(y)+(x)(y^2)=6

OpenStudy (anonymous):

\[x ^{2}y+xy ^{2}=6\]

OpenStudy (anonymous):

with respect to? x or y?

OpenStudy (anonymous):

find: dy/dx

OpenStudy (anonymous):

2xy+y^2 = 0

OpenStudy (anonymous):

how did you do it?

OpenStudy (anonymous):

\[x ^{2}y+xy ^{2}=6\] \[2xy+x^2y'+y^2+2xyy'=0\] solve for \(y'\)

OpenStudy (anonymous):

you are thinking of \(y=f(x),y'=f'(x)\) so you need the product rule for both parts

OpenStudy (anonymous):

\[x ^{2}+2xy+2xy+y ^{2}=0\]

OpenStudy (anonymous):

thats what i got using the product rule

OpenStudy (anonymous):

you also need the chain rule because \(y\) is a function of \(x\) so the derivative of \(y^2\) is \(2yy'\) as the derivative of \(f^2(x)\) is \(2f(x)f'(x)\)

OpenStudy (anonymous):

i still need help guys i dont get it

OpenStudy (anonymous):

do know about product rule?

OpenStudy (anonymous):

yes i know all the rules i just can't figure out this specific problem

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x ^{2}y + x y^{2}) = 0\] now apply product rule

OpenStudy (anonymous):

\[x ^{2}+2xy+2xy+y ^{2}=)\] is that correct?

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

it should be like \[x ^{2}y' + 2xy + y ^{2}+ 2xy' = 0\]

OpenStudy (phi):

notice \[ \frac{d}{dx} y= \frac{dy}{dx} \]

OpenStudy (anonymous):

we can also write \[\frac{ dy }{ dx } = y'\]

OpenStudy (phi):

as compared to \[ \frac{d}{dx} x= \frac{dx}{dx}=1\]

OpenStudy (anonymous):

im so confused i honestly don't get this dx/dy or d/dy stuff

OpenStudy (anonymous):

first you have to know about you are doing derivative to w.r.t to x or y

OpenStudy (phi):

If you need more info on derivatives, these might help http://www.khanacademy.org/math/calculus/differential-calculus/v/calculus--derivatives-1 when you get time. meanwhile, I assume you know how to take the derivative wrt x of \[ \frac{d}{dx} (y=x)\] you get \[ \frac{d}{dx} y=\frac{d}{dx} x\] \[ \frac{dy}{dx} =\frac{dx}{dx} \] \[ \frac{dy}{dx} =1 \] or this problem \[ \frac{d}{dx} (y=x^2)\] you get \[ \frac{dy}{dx} =\frac{d}{dx}x^2 \] \[ \frac{dy}{dx} =2x\frac{dx}{dx} \] \[ \frac{dy}{dx} =2x \]

OpenStudy (phi):

now for a product rule problem d (uv) = u dv + v du \[ \frac{d}{dx} x^2 y = x^2 \frac{d}{dx}y + y \frac{d}{dx}x^2 \] \[ \frac{d}{dx} x^2 y = x^2 \frac{dy}{dx} + y 2x\frac{dx}{dx} \] \[ = x^2 \frac{dy}{dx} + 2xy \]